If i got a 75% on my final and it's worth 10% of my grade and right now i have a 90% in the class what will my final grade be?
Answer by stanbon(75887)

If i got a 75% on my final and it's worth 10% of my grade and right now i have a 90% in the class what will my final grade be?
----
Points out of 100 = 0.75*10 + 0.90*90 = 7.5+81 = 88.5
---
Ans:: 88.5% or 88.5 out of 100
------------------
Cheers,
Stan H.
===================